OpenStudy (anonymous):

How many pairs (A,B) of positive integers are there such that A and B are factors of 6^6 and a is A factor of B

OpenStudy (anonymous):

Well \[ 6^6 = 2^63^6 \]Any factors will be in the form: \[ 2^n3^m \]

OpenStudy (anonymous):

WE have \(n\) and \(m\) range from 0 to 6, so we have 7 choices for n and 7 choices for m.

OpenStudy (anonymous):

So simply considering choices for B, we'd have \(7^2\). But each of them has their own way of getting values of A.

OpenStudy (anonymous):

If we say \[ B = 2^n3^m \]Then the combinations we could expect to have for A would be \((n+1)\cdot (m+1)\).

OpenStudy (anonymous):

So I get \[ \sum_{m=0}^7\sum_{n=0}^7(m+1)(n+1) \]

OpenStudy (anonymous):

Which is \[ \sum_{m=0}^7 (m+1)\sum_{n=0}^7 n +1 = \sum_{m=0}^7 (m+1)\cdot 36 = \sum_{m=0}^7 (m+1)\cdot 36 = 36\cdot 36 = 1296 \]

OpenStudy (anonymous):

Does this answer check out or not?

OpenStudy (anonymous):

for my question.......answer is never to be over 1000 its only 700+

OpenStudy (anonymous):

Hmmm, I guess that \(0\) doesn't count?

OpenStudy (anonymous):

Though, I don't think my answer will have \(0\).

OpenStudy (anonymous):

yea i think (1,6) is counted

OpenStudy (anonymous):

is (1,1) counted?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you understand when I say: \[ B= 2^n3^m \]

OpenStudy (anonymous):

Then we say \[ A = 2^{n'}3^{m'} \]where \(n'\leq n\) and \(m'\leq m\)

OpenStudy (anonymous):

Then we know that A is a factor of B.

OpenStudy (anonymous):

oh wait!! I put 7 instead of 6! So I should get: \[ \sum_{m=0}^6\sum_{n=0}^6(m+1)(n+1) = \sum_{m=0}^6 (m+1)\sum_{n=0}^6 (n+1) = 28\cdot 28 = 784 \]

OpenStudy (anonymous):

???